# Cross-modal Cognitive Consensus guided Audio-Visual Segmentation

### v1

@@ IV-A2 Implementation Details
We implement the proposed method using PyTorch [71]. We use Swin Transformer (Swin) [72] and Pyramid Vision Transformer v2 (PVTv2) [73] as the visual backbones. The channel sizes of the four stages are for Swin and for PVTv2. The visual backbones, including the corresponding classification heads, are pre-trained on ImageNet-1K [67] dataset and the weights of the classification heads are frozen. For audio encoders, the widely adopted VGGish [74] and PANNs [75] pre-trained on AudioSet [68] are used to extract audio features. Since the VGGish does not have a classification head, we trained an additional head for classification using audio features from AudioSet. All of the video frames are resized to without any data augmentation or post-process methods, and the audios are clipped to 1-second splits. The channel size of is set to 256. In all experiments, the models are optimized by Adam [76] and the learning rate is 0.0001. The batch size is set to 4 and the number of training epochs is 20 for the semi-supervised S4 setting and 60 for the fully supervised MS3 setting. We train our model on an RTX 3090 GPU.
@@ IV-B Comparison with State-of-the-Art Methods
Method | Backbone | mIoU () | F-score () | ||
@@
LVS[40] | ResNet-18 2 | 37.9 | 29.5 | 51.0 | 33.0 |
MSSL[39] | ResNet-18+CRNN | 44.9 | 26.1 | 66.3 | 36.3 |
3DC[77] | ResNet-152 (3D) | 57.1 | 36.9 | 75.9 | 50.3 |
SST[24] | ResNet-101 | 66.3 | 42.6 | 80.1 | 57.2 |
iGAN[78] | Swin | 61.6 | 42.9 | 77.8 | 54.4 |
LGVT[79] | Swin | 74.9 | 40.7 | 87.3 | 59.3 |
Baseline[9] | ResNet-50+VGGish | 72.8 | 47.9 | 84.8 | 57.8 |
Baseline[9] | PVTv2+VGGish | 78.7 | 54.0 | 87.9 | 64.5 |
Baseline-Re | ResNet-50+VGGish | 72.7 | 46.6 | 83.2 | 56.3 |
Baseline-Re | PVTv2+VGGish | 78.8 | 54.0 | 87.8 | 65.1 |
C3N (Ours) | Swin+VGGish | 77.5 | 53.5 | 86.9 | 64.7 |
C3N (Ours) | Swin+PANNs | 77.8 | 54.4 | 87.0 | 65.0 |
C3N (Ours) | PVTv2+VGGish | 80.0 | 56.9 | 88.8 | 67.4 |
C3N (Ours) | PVTv2+PANNs | 80.4 | 57.0 | 88.9 | 68.0 |
We use the mIoU and F-score metrics to quantitively compare our method with state-of-the-art methods [40, 39, 77, 24, 78, 79, 9] under the S4 and MS3 settings. We show the performance of the AVS baseline and the authors’ re-implementation of other SSL/VOS/Salient Object Detection (SOD) methods on AVSBench reported in [9]. In addition, we run the authors’ open-source code on our experimental devices under the default settings in [9] to obtain the quantitive performance of the baseline models. The same code running conditions are also employed to get the results of our C3N.
@@
Table I shows the results of the test set. The proposed C3N with PVTv2 [73] as the visual backbone and VGGish [74] or PANNs [75] as the audio encoder achieves remarkable performance and outperforms the baseline method by a large margin. For the simple S4 setup, compared with the baseline with PVTv2 as the visual backbone, the gains achieved by C3N (PVTv2+VGGish) and C3N (PVTv2+PANNs) are 1.2, 1.6 in mIoU and 1.0, 1.1 in F-score, respectively. For the difficult MS3 setting with multiple sounding objects, thanks to the guidance of semantic-level cognitive consensus, C3N (PVTv2+VGGish) and C3N (PVTv2+PANNs) achieve higher performances with a large margin of 2.9, 3.0 mIoU, and 2.3, 2.9 under F-score metric.
@@
We also adopt Swin Transformer [72], a universal visual backbone, to evaluate the general effectiveness of our cognitive consensus-based method. Our C3N with Swin Transformer and VGGish as backbones achieves 77.5 mIoU and 86.9 F-score under the S4 setting and 53.5 mIoU and 64.7 F-score under the MS3 setting. While the results of C3N (Swin+PANNs) are slightly higher. The performance of C3N with Swin as the backbone is close to or even surpasses the PVTv2 version baseline under the MS3 setting, and far higher than the ResNet version under the S4 and MS3 settings.
@@ IV-C1 Analysis of key components
We conduct ablation study experiments with all four backbone combinations. In the experiments, we investigate the effects of audio-visual feature-level interaction, cognitive consensus relies merely on visual modality, and the full audio-visual cognitive consensus. The results are shown in Table II.
@@
The first row shows the performance without audio features and cognitive consensus, i.e. with only video frames as input. Under all four backbone combinations, as in [9], visual-only input does not cause a performance drop by a large margin for the S4 setting while causing a distinct performance drop under the MS3 setting. It indicates that for videos with multiple sounding sources, the introduction of the audio signal and alignment between audio-visual modalities are especially essential to the final prediction. Moreover, in the second row, we introduce the ”CC-V”, which means that the labels with the highest visual semantic confidence are fed into the framework. The results show that only relying on semantic-level guidance from a single modality cannot bring noticeable improvements.
@@
Then we add semantic-level cognitive consensus in the third row. It means the audio-visual semantic-level cognitive consensus is introduced but does not yet include audio features. The audio-visual semantic-level cognitive consensus leads to performance gains for all four groups. In detail, under the MS3 setting, mIoU increases by 1.6, 0.8, 1.5, and 1.9 respectively, and F-score increases between 0.9 to 2.5. For the S4 setting, mIoU increases by around 1 under each backbone combination, and F-score also increases by various points. Furthermore, compared with the results in the fourth row that add audio features, the results in the third row are competitive under the S4 setting. The above results demonstrate that the model performance can be improved through the cross-modal cognitive consensus-based method.
@@
We incorporate audio features and conduct feature-level interaction in the fourth row. The results represent that introducing audio features to the visual input is effective. The reason is that powerful pre-trained audio encoders can capture the dense information of the audio signal. Moreover, the non-local blocks extract cross-modal pixel-wise correlations, which is beneficial to the integration of the two modalities. Nevertheless, such cross-modal interaction is inexplicit due to the absence of semantic-level cognitive consensus, which leads to errors during the segmentation.
@@
Finally, we add both audio features and semantic-level cognitive consensus to the model, i.e. the proposed C3N. It can be noticed that with the cognitive consensus-based cross-modal alignment, the performance of the model improves remarkably under all four backbone settings. In particular, under the MS3 setting, the gains of mIoU are 3.3, 2.8, 2.6, 2.7 respectively and F-score increases by 3.7, 3.6, 2.5 and 3.1. For the S4 setting, mIoU and F-score metrics generally rise by around 1 in all four experimental groups. The above results demonstrate the audio-visual cognitive consensus inference and feedback in our method facilitates explicit cross-modal alignment and improves AVS model performance.
@@ IV-C2 Analysis of Hyperparameters
The two hyperparameters involved in the C3N framework are balance coefficients and in Equation (4). Thue, we conduct multiple settings for the two parameters on the validation set of the S4 and MS3 scenarios under the Swin+PANNs backbone setting. The detailed quantitive results are shown in Table III.
Experimental results show that the mIoU and F-score metrics under different parameter settings (i.e. ==0.10, ==0.25, =0.10 and =0.25, =0.25 and =0.10) fluctuate within 0.5 on the validation set, which is a relatively small range. Finally, we choose ==0.10 as the balance coefficients to construct the confidence re-weighted matrix.
@@ IV-D1 Qualitative analysis
In Fig. 4, we present two AVS examples on the test set of the AVSBench dataset for a qualitative comparison between the baseline method and our C3N. We show segmentation masks of two baseline models (i.e. ResNet-50+VGGish version and PVTv2+VGGish version) and two C3N models (i.e. Swin+VGGish version and PVTv2+VGGish version) with the same audio encoder to ensure objective comparison.
As shown in the left panel, two guitars are simultaneously played. The Res+VGGish version and PVTv2+VGGish version baselines fail to accurately depict both sounding guitars and predict many false positive areas. However, our proposed method uses CCAM to refine the visual features with the guidance of semantic-level cognitive consensus. As a result, our C3N (Swin+VGGish) and C3N (PVTv2+VGGish) segment the sounding object’s contour more precisely.
For the right sample of Fig. 4, the video shows a child playing with a dog as the child continuously makes giggling sounds and the dog is silent. The example is hard because the color of the dog is significantly different from the background, which may mislead the model into incorrectly segmenting the dog as well. The baseline method (Res+VGGish) even completely ignores the sounding child sometimes. The baseline (PVTv2+VGGish) and C3N (Swin+VGGish) segment the dog and child simultaneously. The reason why our C3N (Swin+VGGish) makes false segmentation of the dog is that the Swin generates unbalanced confidence scores and overly attends to the dog in the image, despite the dog and kid should be equally treated without audio guidance, which causes the unified-modal label shifts to the dog. However, the C3N (PVTv2+VGGish) overcomes distraction from the other object and accurately segments the sounding child with tolerable false positive areas.
@@ IV-D2 Effect of cross-modal cognitive consensus inference
The C3IM integrates the classification confidence of distinct modalities with similarities between modality-specific labels to extract the semantic-level unified-modal label. In Fig. 5, we present two visualization examples of the cross-modal cognitive consensus inference process. For ease of illustration, we only show top-5 confidence classes for audio and visual modalities. Note that the audio and visual confidence scores correspond to and in Equation. (4). The highest semantic similarity value, confidence re-weighted value, and the corresponding unified-modal label are framed.
@@
In the bottom example, the sounding object is the violin. This example is more difficult because other prominent targets are present in the visual scene, such as humans and the piano. The sounding violin occupies only a small part of the frame. Column (b) presents the grand piano has the highest confidence and the model incorrectly regard the violin as an oboe. However, after combining it with the audio confidence and semantic similarities, the inferred cognitive consensus still emphasizes the correct violin, fiddle class, which is shown in column (d). According to such guidance information, the model correctly segments the violin. The above visualization results confirm the validity of our C3IM.
@@ IV-D3 Effect of cognitive consensus guided Attention
The Cognitive Consensus guided Attention Module (CCAM) aims to highlight the local feature elements of the sounding object based on inferred cross-modal cognitive consensus. For comparison, we visualize the attention matrices within the audio-visual attention module of the AVS baseline [9] and the spatial attention map within the CCAM of the proposed C3N. In practice, we remove the non-linear activations of CCAM.
@@
In Fig. 6, the heat maps of attention maps of the AVS baseline and the CCAM are shown in the third and fourth rows, respectively. Note that in the fourth row, only the semantic-level alignment of audio and visual modalities is performed and audio features are absent, while the third row shows the audio-visual feature-level attention. In the left example, the violin produces sounds and other objects keep silent. In the third row, although the AVS baseline locates the violin, other irrelevant regions are also highlighted to varying degrees. Nevertheless, in the fourth row, only the sounding violin is emphasized in the spatial attention map of the CCAM. In the right sample, the audio-visual situation is more complex, as two instrument sounds are mixed, and the scene switches in the video. It requires the model to accurately perceive the sounding objects in distinctive scenes. The AVS baseline incorrectly highlights drummers, however, in the fourth row, with the help of cross-modal cognitive consensus, the CCAM heat maps locate the sounding drum and guitar in the first three frames and the last two frames respectively, which is consistent with ground truths.
The above results demonstrate the CCAM allows accurately perceiving and localizing the sounding object only relying on semantic-level cognitive consensus and facilitates the subsequent cross-modal feature fusion and segmentation.

### v4

@@ IV-A2 Implementation Details
We implement the proposed method using PyTorch [96]. We use Swin Transformer (Swin) [97] and Pyramid Vision Transformer v2 (PVTv2) [98] as the visual backbones. The channel sizes of the four stages are for Swin and for PVTv2. The visual backbones, including the corresponding classification heads, are pre-trained on ImageNet-1K [88] dataset and the weights of the classification heads are frozen. For audio encoders, the widely adopted VGGish [99], PANNs [100], and BEATs [101] pre-trained on AudioSet [89] are used to extract audio features. Since the VGGish does not have a classification head, we trained an additional head for classification using audio features from AudioSet. All of the video frames are resized to , and the audio clips are clipped to 1-second splits. Due to the limitation of the data scale of AVSBench, we perform the ColorJitter and RandomHorizonFlip data augmentation strategy. The channel size of is set to 256. In all experiments, the models are optimized by Adam [102] and the initial learning rate is 0.0001. The batch size is set to 8 and the number of training epochs is 20 for the S4 setting. For the MS3 setting, the batchsize is 4 and the number of training epochs is 60.
@@ IV-B Comparison with State-of-the-Art Methods
Method | Backbone | mIoU | F-score | ||
@@
LVS[62] | ResNet-18 2 | 37.94 | 29.45 | 0.510 | 0.330 |
MSSL[61] | ResNet-18+CRNN | 44.89 | 26.13 | 0.663 | 0.363 |
3DC[103] | ResNet-152 (3D) | 57.10 | 36.92 | 0.759 | 0.503 |
SST[47] | ResNet-101 | 66.29 | 42.57 | 0.801 | 0.572 |
iGAN[104] | Swin | 61.59 | 42.89 | 0.778 | 0.544 |
LGVT[105] | Swin | 74.94 | 40.71 | 0.873 | 0.593 |
TPAVI[14] | ResNet-50+VGGish | 72.79 | 47.88 | 0.848 | 0.578 |
TPAVI[14] | PVTv2+VGGish | 78.74 | 54.00 | 0.879 | 0.645 |
TPAVI-Re | ResNet-50+VGGish | 72.69 | 46.56 | 0.832 | 0.563 |
TPAVI-Re | ResNet-50+PANNs | 73.06 | 48.36 | 0.834 | 0.588 |
TPAVI-Re | PVTv2+VGGish | 78.78 | 54.00 | 0.878 | 0.651 |
TPAVI-Re | PVTv2+PANNs | 78.98 | 55.13 | 0.879 | 0.663 |
AVSC[26] | PVTv2+BEATs | 81.29 | 59.50 | 0.886 | 0.657 |
LDM[23] | PVTv2+VGGish | 81.38 | 58.18 | 0.902 | 0.709 |
BG[24] | PVTv2+VGGish | 81.71 | 55.10 | 0.904 | 0.668 |
ECMVAE[25] | PVTv2+VGGish | 81.74 | 57.84 | 0.901 | 0.708 |
ECMVAE-Re | PVTv2+VGGish | 81.64 | 57.44 | 0.899 | 0.690 |
ECMVAE-Re | PVTv2+PANNs | 81.93 | 58.04 | 0.900 | 0.708 |
BAVS[27] | PVTv2+BEATs | 82.68 | 59.63 | 0.898 | 0.659 |
C3N (Ours) | Swin+VGGish | 81.63 | 56.38 | 0.897 | 0.663 |
C3N (Ours) | Swin+PANNs | 81.87 | 57.09 | 0.898 | 0.671 |
C3N (Ours) | Swin+BEATs | 81.95 | 58.14 | 0.900 | 0.683 |
C3N (Ours) | PVTv2+VGGish | 82.67 | 60.11 | 0.904 | 0.697 |
C3N (Ours) | PVTv2+PANNs | 82.94 | 61.67 | 0.906 | 0.713 |
C3N (Ours) | PVTv2+BEATs | 83.11 | 61.72 | 0.908 | 0.722 |
We use the mIoU and F-score metrics to quantitively compare our method with state-of-the-art methods [62, 61, 103, 47, 104, 105, 14, 26, 23, 24, 25, 27] under the S4 and MS3 settings. We show the performance of SSL/VOS/Salient Object Detection (SOD) methods on AVSBench reported in [14] in the first panel of Table I. In addition, in the second panel, we also show the performance of many brand-new AVS methods for comprehensive comparison, and we re-implement the official source code of TPAVI [14] and ECMVAE [25]. In addition, we also re-implement them with PANNs [100] as the audio backbone for a fair assessment. Finally, in the third panel, we show the performance of our C3N under six different backbone combinations.
@@
Table I shows the results of the test set. The proposed C3N with PVTv2 [98] as the visual backbone and VGGish [99] or PANNs [100] as the audio encoder achieves remarkable performance and outperforms other AVS methods. For the simple S4 setup, compared with BAVS [27] with PVTv2 as the visual backbone and strong BEATs [101] as the audio backbone, our C3N (PVTv2+PANNs) gains higher mIoU and F-score and C3N (PVTv2+VGGish) achieves comparable performance. For the difficult MS3 setting with multiple sounding objects, thanks to the guidance of semantic-level cognitive consensus, C3N (PVTv2+VGGish) and C3N (PVTv2+PANNs) achieve higher performances than BAVS by 0.48, 2.04 in mIoU and 0.038, 0.054 in F-score, respectively.
@@
We also adopt Swin Transformer [97] to evaluate the general effectiveness of our cognitive consensus-based method. Our C3N (Swin+VGGish) achieves 81.63 mIoU and 0.897 F-score under the S4 setting and 56.38 mIoU and 0.663 F-score under the MS3 setting. While the results of C3N (Swin+PANNs) are higher. Despite the PVTv2 backbone is stronger than the Swin backbone, the performance of C3N with Swin as the backbone is comparable with some PVTv2-based AVS methods and far higher than the classical TPAVI (ResNet version) under the S4 and MS3 settings. In addition, we also adopt BEATs [101] as the audio backbone, which leads to mIoU and F-score performance improvement compared with C3N with the VGGish or PANNs as the audio backbone under the S4 and MS3 settings. Specifically, our C3N (Swin+BEATs) outperforms C3N (Swin+PANNs) by 1.05 in mIoU and 0.012 in F-score under the MS3 setting and achieves slightly higher mIoU and F-score under the S4 setting. C3N (PVTv2+BEATs) achieves higher performance than C3N (PVTv2+PANNs) by 0.05 in mIoU and 0.009 in F-score under the MS3 setting, and 0.17 in mIoU and 0.002 in F-score under the S4 setting.
@@ IV-C1 Analysis of key components
In Table II, we conduct ablation experiments under all four backbone combinations. In the experiments, we investigate the effects of audio-visual feature-level interaction, cognitive consensus relies merely on visual modality, and the full audio-visual cognitive consensus.
@@
The first row shows the performance without audio features and cognitive consensus, i.e. with only video frames as input. Under all four backbone combinations, as in [14], visual-only input does not lead to a performance drop by a large margin for the S4 setting while causing a distinct performance drop under the MS3 setting. It indicates that for videos with multiple sounding sources, the introduction of the audio signal and alignment between audio-visual modalities are especially essential to the final prediction. Moreover, in the second row, we introduce the “CC-V”, which means that the labels with the highest visual semantic confidence are fed into the framework. The results show that the performance after the introduction of “CC-V” is improved. However, relying only on semantic-level guidance from a single modality brings limited improvement.
@@
Then we add semantic-level cognitive consensus in the third row. It means the audio-visual semantic-level cognitive consensus is introduced but does not yet include audio features. The audio-visual semantic-level cognitive consensus leads to performance gains for all four groups. In detail, under the MS3 setting, mIoU increases by 1.51, 2.91, 3.55, and 1.90 respectively, and F-score increases between 0.026 to 0.034. For the S4 setting, mIoU increases by around 1 under each backbone combination, and the F-score also increases by various points. Furthermore, compared with the results in the fourth row that add audio features, the results in the third row are competitive under the S4 setting. The above results demonstrate that the model performance can be improved through the cross-modal cognitive consensus-based method.
@@
We incorporate audio features and conduct feature-level interaction in the fourth row. The results represent that introducing audio features to the visual input is effective. The reason is that powerful pre-trained audio encoders can capture the dense information of the audio signal. Moreover, the non-local blocks extract comprehensive cross-modal pixel-wise correlations, which is beneficial to the integration of the two modalities. Nevertheless, such cross-modal interaction is inexplicit due to the absence of semantic-level cognitive consensus, which leads to some errors during the segmentation.
@@
Finally, we add both audio features and semantic-level cognitive consensus to the model, i.e. the proposed C3N. It can be noticed that with the cognitive consensus-based cross-modal alignment, the performance of the model improves remarkably under all four backbone settings. In particular, under the MS3 setting, the gains of mIoU are 3.33, 2.75, 3.38, 2.91 respectively and F-score increases by 0.038, 0.022, 0.034 and 0.026. For the S4 setting, mIoU and F-score metrics generally rise by around 1 and 0.01 in all four experimental groups, respectively. The above results demonstrate the audio-visual cognitive consensus inference and feedback facilitates explicit cross-modal alignment and improves AVS model performance.
@@ IV-C2 Analysis of Hyperparameters
The two hyperparameters involved in the C3N framework are balance coefficients and in Equation (6). Thus, we conduct multiple settings for the two parameters on the validation set of the S4 and MS3 scenarios under the Swin+PANNs backbone setting. The detailed quantitive results are shown in Table III. Experimental results show that the F-score metrics under different settings fluctuate within a relatively small range. Furthermore, the model performance is the overall best when ==0.10.
IV-C3 Analysis of Model Complexity
Complexity | Baseline | C3N | () |
FLOPs (GB) | 161.36 | 161.39 | 0.019 |
#Params (MB) | 171.46 | 177.15 | 3.32 |
In Table IV, we evaluate the model complexity of the baseline and C3N methods under the same backbone setting, where the baseline means removing the proposed C3IM and CCAM from the C3N model. means the increased FLOPs/Params of our C3N as a proportion of the baseline model. The inference FLOPs and the number of parameters of the C3N increase by 0.03GB and 5.69 MB, which account for 0.019 and 3.32 of the baseline, respectively. Considering the improvement of the model performance, the increased model complexity is tolerable.
@@ IV-D1 Qualitative analysis
In Fig. 5, we present two AVS examples on the validation and test set of the AVSBench dataset for a qualitative comparison between the baseline method TPAVI [14] and our C3N. We show segmentation masks of two TPAVI models (i.e. ResNet-50+VGGish version and PVTv2+VGGish version) and two C3N models (i.e. Swin+VGGish version and PVTv2+VGGish version).
As shown in the left panel, there is a sounding ambulance with an alarm lamp and a normal van in the image. The Res+VGGish and PVTv2+VGGish version TPAVI models either fail to segment the sounding ambulance or incorrectly segment the normal van. The reason is that the dense feature-level interactions of the TPAVI model lead to the mislocalization of the representationally similar but semantically different object. However, our C3N utilizes cognitive consensus to align the audio and visual modalities from the semantic level and make correct localization and precise segmentation.
For the right sample of Fig. 5, the video shows a child playing with a dog as the child continuously makes giggling sounds and the dog is silent. The example is hard because the color of the dog is significantly different from the background, which may mislead the model into incorrectly segmenting the dog as well. The TPAVI model (Res+VGGish) even completely ignores the sounding child sometimes. The TPAVI (PVTv2+VGGish) and C3N (Swin+VGGish) segment the dog and child simultaneously. The reason why our C3N (Swin+VGGish) makes false segmentation of the dog is that the Swin generates unbalanced confidence scores and overly attends to the dog in the image, despite the dog and kid should be equally treated without audio guidance, which causes the unified-modal label shifts to the dog. However, the C3N (PVTv2+VGGish) overcomes distraction from the other object and accurately segments the sounding child.
@@ IV-D2 Effect of cross-modal cognitive consensus inference
The C3IM integrates the audio and visual classification confidence scores and label embedding similarities to extract cognitive consensus. In Fig. 6, we present two examples of the cross-modal cognitive consensus inference process. For ease of illustration, we only show top-5 confidence classes for audio and visual modalities. Note that the audio and visual confidence scores correspond to and in Equation. (6). The highest semantic similarity value, confidence re-weighted value, and the corresponding unified-modal label are framed.
@@
In the bottom example, the sounding object is the violin. This example is more difficult because other prominent targets are present in the visual scene, such as humans and the piano. The sounding violin occupies only a small part of the frame. Column (b) presents the grand piano has the highest confidence and the model incorrectly regards the violin as an oboe. However, after combining it with the audio confidence and semantic similarities, the inferred cognitive consensus still emphasizes the correct violin, fiddle class, which is shown in column (d). According to such guidance information, the model correctly segments the violin. The above visualization results confirm the validity of our C3IM.
@@ IV-D3 Effect of cognitive consensus guided attention
The Cognitive Consensus guided Attention Module (CCAM) aims to highlight the local feature elements of the sounding object based on inferred cross-modal cognitive consensus. For comparison, we visualize the attention matrices within the audio-visual attention module of TPAVI [14] and the spatial attention map within the CCAM of the proposed C3N. In practice, we remove the non-linear activations of CCAM.
@@
In Fig. 7, the heat maps of attention maps of the TPAVI baseline and the CCAM are shown in the third and fourth rows, respectively. Note that in the fourth row, only the semantic-level alignment of audio and visual modalities is performed, while the third row shows the audio-visual feature-level attention. In the left example, the violin produces sounds and other objects keep silent. In the third row, although the TPAVI baseline locates the violin, other irrelevant regions are also highlighted to varying degrees. Nevertheless, in the fourth row, only the sounding violin is emphasized in the spatial attention map of the CCAM. In the right sample, the situation is more complex, as two instrument sounds are mixed, and the scene switches in the video. It requires the model to accurately locate the sounding objects in distinctive scenes. The TPAVI baseline incorrectly highlights drummers. However, with the help of cognitive consensus, the CCAM heat maps locate the sounding drum and guitar in the first three frames and the last two frames respectively, which is consistent with ground truths. The above results demonstrate the CCAM allows accurately localizing the sounding object only relying on semantic-level cognitive consensus.
IV-D4 Effect of cross-modal feature fusion
The role of the cross-modal Non-local block-based feature fusion is to perform dense audio-visual feature-level interaction to calibrate the visual features via the discriminative audio features, which enforces highlighting the visual context of sounding objects to distinguish the sounding and silent objects with similar semantics. In Fig. 8, we show the visual feature heatmaps obtained by averaging the feature and the segmentation predictions output by our C3N with or without the cross-modal Non-local blocks. In the first row, there are two men with the same semantic label in a frame while one of them is speaking. The C3N without the cross-modal Non-local blocks incorrectly highlights and segments both of them. Our C3N with the cross-modal Non-local blocks identifies the speaking man by calibrating the visual features via the discriminative audio features to make correct segmentation. In the second row, there are two pianos, one of which is being played. The C3N without the cross-modal Non-local blocks incorrectly highlights and segments the silent piano. Whereas the C3N with the cross-modal Non-local blocks highlights the piano being played by hands and predicts the accurate segmentation mask.
IV-D5 Failure cases
We visualize three failure cases in Fig. 9. In the first row, the failure example is caused by highly similar visual appearances with identical semantics. Specifically, the semantics and appearances of the two guitars are highly similar and both have a hand on them, and it is difficult for the model to tell which guitar is sounding. In the second row, the failure example is caused by object occlusion. Specifically, the barking dog is covered by the man in a large area, which causes the C3N model to mislocate and output failure prediction. In the third row, the volume of the input audio is so low that the audio features lose the discriminability towards the sounding object, so the model cannot localize the sounding violin and just outputs the mask with several tiny positive areas. In the above cases, it is difficult to identify and segment the sounding object based on the limited information of a single frame. In the future, we will concentrate on mining the inter-frame temporal context compensational information of the audio-visual inputs to address problems in these cases.
IV-D6 Evaluation under different genres of videos
ambulance siren | baby laughter | gun shooting | cat meowing | chainsawing trees | coyote howling | dog barking | driving buses | ||||||||
mIoU | F-score | mIoU | F-score | mIoU | F-score | mIoU | F-score | mIoU | F-score | mIoU | F-score | mIoU | F-score | mIoU | F-score |
79.81 | 0.863 | 82.62 | 0.894 | 72.57 | 0.851 | 87.77 | 0.934 | 65.45 | 0.789 | 83.36 | 0.914 | 87.26 | 0.927 | 84.08 | 0.891 |
female singing | helicopter | horse clip-clop | lawn mowing | lions roaring | male speech | bird singing | playing guitar | ||||||||
mIoU | F-score | mIoU | F-score | mIoU | F-score | mIoU | F-score | mIoU | F-score | mIoU | F-score | mIoU | F-score | mIoU | F-score |
81.70 | 0.878 | 76.96 | 0.880 | 81.16 | 0.895 | 86.72 | 0.931 | 91.13 | 0.959 | 90.61 | 0.958 | 87.96 | 0.949 | 87.71 | 0.946 |
playing glockenspiel | playing piano | playing tabla | playing ukulele | playing violin | race car | typing keyboard | Class Average | ||||||||
mIoU | F-score | mIoU | F-score | mIoU | F-score | mIoU | F-score | mIoU | F-score | mIoU | F-score | mIoU | F-score | mIoU | F-score |
83.45 | 0.934 | 87.30 | 0.940 | 87.07 | 0.920 | 76.59 | 0.873 | 74.35 | 0.861 | 85.74 | 0.924 | 88.74 | 0.949 | 83.056.21 | 0.9070.041 |
To evaluate the effectiveness of our C3N under different genres of videos as inputs, we conduct the per-category performance test on 23 categories for our C3N (PVTv2+PANNs) under the S4 setting, which has the video category annotations, and the average and standard deviation of mIoU and F-score metrics across all categories are also reported. As shown in Table V, our C3N (PVTv2+PANNs) achieves an average mIoU of 83.05 and an average F-score of 0.907 across all 23 categories, with their respective standard deviation of 6.21 and 0.04, which demonstrates that the performance of our C3N fluctuates within a relatively small range with video category changes.
IV-D7 Evaluation under varying sound qualities
SNR (dB) | mIoU | F-score | ||
S4 | MS3 | S4 | MS3 | |
No noise | 82.94 | 61.67 | 0.906 | 0.713 |
20dB | 82.86 | 61.52 | 0.905 | 0.713 |
10dB | 82.79 | 60.60 | 0.904 | 0.703 |
0dB | 82.65 | 59.16 | 0.903 | 0.693 |
-10dB | 82.27 | 55.23 | 0.900 | 0.674 |
-20dB | 81.50 | 51.08 | 0.896 | 0.645 |
To evaluate the adaptivity of our method under varying sound qualities, we add Gaussian white noise of a variety of signal-to-noise ratios (SNR) to the input audio to simulate different quality audio signals. As demonstrated in Table VI, the segmentation performance of our C3N declines with the SNR decreasing. Under the S4 setting, since there is only one sounding object in the video, our C3N is slightly influenced by the added noise, i.e. the mIoU and F-score decrease in the range of 1.5 mIoU and 0.01 F-score with 20dB to -20dB SNR audio inputs, respectively. Due to the presence of multiple sounding objects in a video under the MS3 setting, our C3N is more sensitive to audio quality compared to the S4 setting. When the SNR decreases from 20dB to 0dB, the performance of our C3N drops less than 3 mIoU and 0.02 F-score compared with C3N with clear audio inputs. Although the performance of C3N significantly degrades when the added noise power is greater than the clear audio (i.e. SNR0dB), our C3N still achieves 55.23 mIoU and 0.674 F-score with -10dB SNR audio, and 51.08 mIoU and 0.645 F-score with -20dB SNR audio. The above results demonstrate the robustness of the proposed C3N under varying sound qualities.
IV-D8 Evaluation under varying visual qualities
SNR (dB) | mIoU | F-score | ||
S4 | MS3 | S4 | MS3 | |
No noise | 82.94 | 61.67 | 0.906 | 0.713 |
30dB | 82.58 | 61.59 | 0.904 | 0.712 |
20dB | 80.81 | 59.40 | 0.894 | 0.697 |
10dB | 74.06 | 52.09 | 0.849 | 0.629 |
To evaluate the adaptivity of our method under noisy visual environments, we add Gaussian noise of varying signal-to-noise ratios (SNR) to the input video frames to simulate different quality videos. In Fig. 10, we show video frames with the noise of different signal-to-noise ratios (SNR) and their respective segmentation prediction output by our C3N (PVTv2+PANNs). When the SNR of the input frame is 30dB or 20dB, our C3N outputs accurate segmentation masks. However, when SNR is decreased to 10dB, the frame is severely distorted, and the quality of the predicted mask is significantly lower compared to the clear frame input. As shown in Table VII, the C3N performance declines with the SNR decreasing. When the SNR drops to 20 dB, the mIoU and F-score slightly reduce by 2.13 and 0.012 under the S4 setting and 2.27 and 0.016 under the MS3 setting. Despite the performance significantly drops when SNR is decreased to 10dB, our C3N still achieves 74.06 mIoU and 0.849 F-score under the S4 setting, and 52.09 mIoU and 0.629 F-score under the MS3 setting. The above experimental results demonstrate the noise robustness of our C3N when encountering not severe visual noise.